use cramers rule or to determine that the system is inconsistent or contains dependent equations:
y=-4x+2
2x=3y+8

Answer by nabla(475)

Put the equations in consistent form:
4x+y=2
2x-3y=8
This gives us a 2x2 system with matrix
4 1 2
2 -3 8
So by Cramer's rule,
x=
2 1
8 -3
----
4 1
2 -3
and
y=
4 2
2 8
----
4 1
2 -3
So, x= (-6-8)/(-12-2)=1
and y= (32-4)/(-14)=-2
Thus we have a unique solution.